Dictionary.com
Thesaurus.com
Showing results for heterozygote. Search instead for merozygote.

heterozygote

American  
[het-er-uh-zahy-goht, -zig-oht] / ˌhɛt ər əˈzaɪ goʊt, -ˈzɪg oʊt /

noun

Genetics.
  1. a person, animal, or other organism with a pair of differing alleles, one dominant and one recessive, of a particular gene.


heterozygote British  
/ ˌhɛtərəʊˈzaɪɡəʊt, -ˈzɪɡəʊt /

noun

  1. an animal or plant that is heterozygous; a hybrid Compare homozygote

"Collins English Dictionary — Complete & Unabridged" 2012 Digital Edition © William Collins Sons & Co. Ltd. 1979, 1986 © HarperCollins Publishers 1998, 2000, 2003, 2005, 2006, 2007, 2009, 2012

Other Word Forms


Etymology

Origin of heterozygote

First recorded in 1900–05; hetero- + zygote


Explanation

A heterozygote is a living thing that has two different versions of a given gene. If a cat inherits a gene for black fur from its mother and one for gray fur from its father, it is a heterozygote. The different forms of a gene are called alleles. Organisms inherit one allele for each gene from their biological parents. If those genes are different — e.g., one for gray fur, and one for black fur — the offspring is a heterozygote with respect to fur color. What color fur will a heterozygous cat have? It depends on which gene is dominant. Since the gene for black fur is dominant, the cat would be black. However, the cat could pass on the non-dominant, or recessive, gene to its offspring.

Keep Reading on Vocabulary.com

Example Sentences

Examples are provided to illustrate real-world usage of words in context. Any opinions expressed do not reflect the views of Dictionary.com.

See Examples For:

Figure 12.5 A test cross can be performed to determine whether an organism expressing a dominant trait is a homozygote or a heterozygote.

From Textbooks ● Jun. 9, 2022

This may occur when the mutant allele somehow interferes with the genetic message so that even a heterozygote with one wild-type allele copy expresses the mutant phenotype.

From Textbooks ● Jun. 9, 2022

The probability of the homozygote or the heterozygote is 1/4 + 1/2 = 3/4 using the sum rule.

From Textbooks ● Jun. 9, 2022

Coat color depends on the A gene, which shows dominance, with the recessive homozygote showing a different phenotype than the heterozygote or dominant homozygote.

From Textbooks ● Apr. 25, 2013

The heterozygote individual in typical cases shows the dominant character.

From Hormones and Heredity by J. T. Cunningham

Random Word

Spin the wheel and learn a new word!

Get a word

Vocabulary.com logo
by dictionary.com

Just like flashcards, but way more fun

Don't take our word for it. Leave the boring studying behind and join VocabTrainer today. You'll learn even faster and never look back.

Take me to Vocabulary.com